### v1

@@ Abstract
TikTok is now a massive platform, and has a deep impact on global events. But for all the preliminary studies done on it, there are still issues with determining fundamental characteristics of the platform. We develop a method to extract a representative sample from a specific time range on TikTok, and use it to collect >99% of posts from a full hour on the platform, alongside a dataset of >99% of posts from a single minute from each hour of a day. Through this, we obtain post metadata, video media data, and comments from a close to complete slice of TikTok. Using this dataset, we report the critical statistics of the platform, notably estimating a total of 117 million posts produced on the day we looked at on TikTok.
@@ 1 Introduction
Data collection methods on TikTok are limited. Search mechanisms on the site are limited to hashtag based search which stops at 1000 returns, fuzzy search, and user and song based search [13]. All of these methods limit the extent to which a true complete slice of the platform can be examined. While TikTok has been offering a Researcher API to selected scholars [14], it has been shown recently, that this API shows significant deviations in terms of metadata and availability of videos when compared to the TikTok website [15]. Additional limitations of the API include restrictive user agreements [16], and the inability to return videos originating from Canada  [^1] https://developers.tiktok.com/doc/research-api-codebook/. We must obtain stronger sampling methods if we are to gain better insight into a platform of TikTok’s magnitude [17]. Without a representative sample from such a method, researchers cannot properly contextualize or validate findings about behaviour on TikTok.
To address this need, we developed a novel method for collecting >99% uploaded to TikTok within a specific timespan, enabling a collection of a close to complete slice of TikTok for the first time.
@@
Considering the focus of research on TikTok and social media platforms in general, we will provide insight on the following points:
@@
As part of this work, we will release the fundamental statistics necessary to understand TikTok, and calibrate studies done on it. This data is critical for grounding findings from TikTok. With that in mind, we will find these statistics with a methodology detailed in this work, and release them at URL.
With this study, we have violated TikTok’s Terms of Service  [^2] https://www.tiktok.com/legal/page/us/terms-of-service/en, specifically the following section: ‘use automated scripts to collect information from or otherwise interact with the Services;’ However, we strongly believe that our work is critical to better understand a platform that is used by billions of users worldwide. Greater platform transparency will lead to higher quality and more trustworthy research, which in turn, will help the community and TikTok work together to help the platform flourish.
@@ 2 Results
We developed a method that, to the best of our knowledge, allows us to collect >99% of the post content on TikTok across the world. We used that method to collect every video made between 5-6pm UTC on the 10th April, 2024 (an arbitrary, mundane date, to obtain baseline data from the platform), and all videos from the 42nd minute of each hour from 10pm UTC on the 9th April, till 10pm on the 10th April, to obtain across day representative data. Here, we use that data to extract what we consider to be the critical statistics of note on TikTok, that will let the research community contextualize findings they have made on TikTok.
@@ 2.1 Statistics
(a) Histogram of comment count.
(b) Histogram of view count.
(c) Histogram of like count.
(d) Histogram of share count.
(e) Histogram of video duration.
Figure 1: Histograms of various statistics in the video metadata, across the 1 hour dataset.
@@ 2.1.1 Estimating overall post volume
For the observed hour on TikTok, we collected $5,157,488$ videos with an average of $1,432.6$ videos [$1,166-1,758$] per second (see Fig. 2(a)). The aggregated numbers per minute range from $80,292$ to $95,915$ (avg. $85,958.1$) videos per minute (Fig. 2(b)).
In order to estimate the videos per hour for an entire day, we collected the videos of the $42^{nd}$ minute of every hour. For our collected hour, the $42^{nd}$ minute is slightly less active ($\sim 96.5\%$) than the average of the other $59$ minutes. Consequently, we used this correction factor to calculate the estimated number of videos per hour. The result can be seen in Fig. 2(c) and shows an interesting two-peaked temporal pattern. These estimated hours aggregate to 116.7 million video posts over a 24-hour time period.
@@
We note that the time distribution of videos per hour looks markedly different from Corso et al. [14]. We also see that there is a 14% higher posting volume at the 0th second of each minute (where the distribution is otherwise flat), implying bot activity, or scheduled posts. View counts for videos posted at the zeroth second get significantly more views (t-test 3.97, p-value: 7e-5), with an average of 3899 views at the zeroth second vs. 2502 views for all other seconds. Videos at the zeroth second more frequently use hashtags implying the desire to reach a broader audience, such as fyp, implying that these posters are using post scheduling and hashtag use to attempt to reach a broader audience.
Understanding the total size of the platform gives much needed context when we want to understand samples of it.
@@ 2.1.2 Understanding global usage
Figure 3: Choropleth of video posting per capita for all countries, determined via the locationCreated tag available on the video post metadata. We normalize by the population of each country to show posts per capita. The counts are corrected for geographical regions that we estimate that we underrepresent.
Figure 4: Regression of country population against video count in this dataset, with the 6 countries furthest from the trend line labelled, as determined by highest and lowest residuals from the linear regression.
We then looked at global usage of TikTok. Research on TikTok frequently looks at country specific slices [22], so understanding actual country level usage is critical for context [20]. We infer post country creation from the locationCreated tag that refers to video origin country  [^3] https://developers.tiktok.com/doc/research-api-specs-query-videos. While this indicator of country origin is likely confounded by VPN usage and other factors, it represents our current best method for assessing post country origin. Our method has slight variation in regional representativeness, so we adjust counts based on known statistics from our method, detailed in Section 4.3.
@@
Using the locationCreated attribute, we can also infer the distribution of when TikTok posts are created in the local time, see Figure 5. In the case of countries spanning multiple timezones, we simply select the first timezone associated with a country from the Olson timezone database  [^5] https://github.com/stub42/pytz/blob/master/tz/zone.tab. Inaccuracies are inevitable due to lack of sub-country location information, and countries spanning multiple timezones, but the general trends should be preserved nonetheless.
@@ 2.2 Video Media
TikTok is a video based platform, and any understanding of the nature of the platform as a whole needs to come with a look at the actual video media data of the platform.
For each post in our dataset that has associated video media data (i.e. not a story-post), we downloaded the video media data. We wanted to gain a view into what kind of topics were discussed on the platform. We therefore used a combination of topic modelling and vision language models to cluster the videos, and describe each cluster. We plot the posts and clusters in Figure 6.
@@
We can see the broad array of topics with the largest being videos of a person talking directly to the camera, along with common activities in everyday life such as sports, transport, and lifestyle videos. In smaller clusters we find more uncommon topics, such as ‘Political Commentary on Police Operations in Pakistan’ (0.025% of videos), ‘Camouflaged Military Soldier Image’ (0.06%), and ‘Palestinian Flag Protest Imagery’ (0.008%).
Now, when we look at specific activity on TikTok, such as political, military, and healthcare content, we can start to understand what proportion of the whole platform it represents.
@@ 2.2.1 Presence of Children
The heavy use of TikTok by children and teenagers has caused concern for safety on the platform. But the lack of a representative sample of TikTok has meant so far we haven’t been able to estimate the true magnitude of child activity on the platform. We know there is a high prevalence of children on the platform, with 50% of 5-7 year olds, and 80% of 12-15 year olds in the UK using the platform [21], but we don’t what percentage of content on the platform features children. We aim to address that here.
We trained a classifier to detect if a video contains a child, with a mean F1 of 0.86 from 5-fold cross-validation on our train/testset (precision 0.88, recall 0.81). We used this classifier to find that 18.13% of the videos in our 24 hour dataset contain children (95% confidence interval: [18.06%, 18.21%] from bootstrapped confidence intervals). This is a strikingly high statistic, and we believe there is not a specific topic for children in our topic model purely because there they appear so commonly and diversely in the videos. We then wanted to know how this measure breaks down by country.
We removed countries where there were insufficient post counts to make a conclusive estimate of the share of videos containing children, based on a binomial test (p=0.05).
@@
Figure 8: Scatter plot of countries, for each country showing the share of the population under the age of 14 against the share of videos from that country containing one or more children, as determined by our video classifier. We have removed countries where there was insufficient data to make a conclusive estimate of the share of videos containing children, based on a binomial test (p=0.05). We label the top 2 and bottom 2 countries with the highest residual from the linear regression, effectively showing countries with the most overrepresented and underrepresented children in TikTok posts, respectively.
Figure 7 shows a choropleth of share of videos containing one or more children per country, adjusted by the share of the population under the age of 14  [^6] https://population.un.org/wpp/ (to avoid showing simply a map of countries with a youthful population). This effectively shows countries where there is a culture of posting videos containing children. Figure 8 shows a scatter plot of countries comparing the share of the population under 14, and the share of videos containing one or more children, including a linear regression. We label the top 3 and bottom 3 countries with the highest residual from the linear regression, effectively showing countries with the most overrepresented and underrepresented children in TikTok posts, respectively. This data therefore lets us couch our understanding of the prominence of children on TikTok in the actual fraction of the platform they take up.
@@ 2.3 Comments
Comments are the richest form of mass engagement data on TikTok, providing language, semantic, and temporal data. We can use them to build understanding of TikTok’s audience.
We collected 18.4 million comments directed at the videos (excluding comments on comments), associated with 13.5 million commenters. We collected these comments in September 2024, 5 months after the day the videos were posted, using Ensemble Data  [^7] https://ensembledata.com/.
@@
When looking at the time distribution of comment creation time, we see a flat distribution (std dev. across all seconds: 652, std dev. across all minutes: 3.7k, with a mean of 306k), indicating minimal scheduled posting of comments, and if automated comment posting is occurring, timing of posting is more sophisticated.
We also looked at average offset between video post and comment time, as shown in Figure 9. We see that for all commenters, the modal time to comment is just under 1 day, and that more frequent commenters comment have a modal peak of 1 hour from post time. We can also see that around 2/3 months after a video is posted, the commenting frequency on the video rapidly declines. This data shows us that while distribution processes on TikTok are fast, users are not engaging with the content until at least a few minutes past post time.
Looking at comment language, 47% of comments are unknown (likely due to the high prevalence of emoji only comments on the platform), 15% are in English, 9% are in Spanish, and 5% are in Arabic. The high percentage of emoji only commenting on TikTok could imply large quantities of inauthentic activity, as emojis would be an easy way to boost a video’s engagement while requiring limited effort on behalf of the bot.
@@ 2.4 Error Analysis
A key feature of our collection method means that we receive a great number of error messages when we attempt to collect posts that either don’t exist, or are not available for other reasons. As to our knowledge this has not previously been reported on, we report on it here.
We received several different types of error message for a requested video, and we report the most common error types, excluding the error message ‘item doesn’t exist’, in Table 2. We can see that an extremely large number of video posts are deleted, or privatized (status_self_see). We also gain an insight into what seems to be the TikTok moderation process, with  7% of videos unavailable at the time of collection due to a review stage (status_reviewing),  1% of videos hidden after audit (status_audit_not_pass), and a small number potentially currently undergoing content classification for moderation (content_classification). We also see two error messages potentially related to geo-location, with cross_border_violation likely related to location based content restrictions, one of which was deployed around the time of the invasion of Ukraine [27].
@@
status_deleted 3,763,773 20.96
status_self_see 3,732,155 20.79
status_reviewing 1,201,251 6.69
status_audit_not_pass 191,690 1.07
content_classification 26,654 0.15
@@
Table 2: Distribution of most common post error messages, excluding ‘item doesn’t exist’. Reported percentage is out of all videos that at some point existed in our combined datasets. So from this we can infer that around 20% of videos on TikTok are deleted.
When we look at deletion rates over different parts of our collection period, we see that of the posts that exist, at 26 days after post time, deleted videos accounted for 12% of videos, and at 215 days after post time, they accounted for 20% of videos.
This data gives us a first insight into inaccessible content on TikTok, and future work can exploit this to bring to light the moderation processes used by the platform, the rate of deletion, and the nature of cross-border content restrictions.

### v3

@@ Abstract
TikTok is now a massive platform, and has a deep impact on global events. But for all the preliminary studies done on it, there are still issues with determining fundamental characteristics of the platform. We develop a method to extract a representative sample from a specific time range on TikTok, and use it to collect >99% of posts from a full hour on the platform, alongside a dataset of >99% of posts from a single minute from each hour of a day. Through this, we obtain post metadata, video media data, and comments from a close to complete slice of TikTok. Using this dataset, we report the critical statistics of the platform, notably estimating a total of 269 million posts produced on the day we looked at on TikTok, that 18% of videos on the platform feature children, and that $\sim$176 years of videos are uploaded every day.
@@ 1 Introduction
Data collection methods on TikTok are limited. Search mechanisms on the site are limited to hashtag based search which stops at 1000 returns, fuzzy search, and user- and song-based search [13]. All of these methods limit the extent to which a true complete slice of the platform can be examined. While TikTok offers a Researcher API to selected scholars [14], it has been established recently that this API shows significant deviations in terms of metadata and availability of videos when compared to the TikTok website [15]. Additional limitations of the API include restrictive user agreements [16] and the inability to return videos originating from Canada  [^1] https://developers.tiktok.com/doc/research-api-codebook/. We must obtain stronger sampling methods if we are to gain contextualized insight into a platform of TikTok’s magnitude [17]. Without representative samples, researchers cannot properly contextualize or validate findings about behaviour on TikTok.
To address this need, we developed a novel method for collecting >99% uploaded to TikTok within a specific timespan, enabling a collection of a close-to-complete slice of TikTok for the first time.
@@
We used the representative slice of TikTok that we collected to examine a number of fundamental phenomena on TikTok. Considering prevailing research themes around TikTok and other social media platforms, we provide insight on the following points:
@@
As part of this work, we also release a collection of fundamental statistics around these phenomena. We expect these will be useful to a wide-array of future studies: both for the studies themselves and to calibrate their findings against the broader backdrop of TikTok behavior.
In collecting the posts and media content for this study, we have violated TikTok’s Terms of Service  [^2] https://www.tiktok.com/legal/page/us/terms-of-service/en, specifically the following section: ‘use automated scripts to collect information from or otherwise interact with the Services;’ However, we strongly believe that our work is critical to better understand a platform that is used by billions of users worldwide. Greater platform transparency will lead to higher quality and more trustworthy research, which in turn, will help the community and TikTok work together to help the platform flourish.
@@ 2 Results
We developed a method that, to the best of our knowledge, allows us to collect >99% of the post content on TikTok across the world. We used that method to collect every post made between 5-6pm UTC on the 10th of April, 2024 (an arbitrary, mundane date, to obtain baseline data from the platform), and all posts from the 42nd minute of each hour from 10pm UTC on the 9th of April, till 10pm on the 10th of April, to obtain an across-day representative data sample. After collection, we collected the video media content and comments of the posts where available. Here, we use that data to extract what we consider to be the critical statistics of note on TikTok, that will let the research community contextualize findings they report on TikTok.
@@ 2.1.1 Estimating overall post volume
For the observed hour on TikTok, we collected $5,157,488$ posts. By looking at error messages for other requested videos, we can determine that there were originally $12,044,235$ posts created in this hour (with many deleted, made private, or otherwise made inaccessible since post time, see Section 2.4), with an average of $3,345.6$ videos per second (minimum-maximum $2,773-9,042$, see Fig. 1(a)). The aggregated numbers per minute range from $189,620$ to $256,743$ (avg. $200,737.3$) videos per minute (Fig. 1(b)). Here, because we want to consider all posts that were available in the original time-span, included no longer available posts, we use the create time that can be extracted from the post ID, see Section 4.1.
In order to estimate the videos per hour for an entire day, we collected the videos of the $42^{nd}$ minute of every hour. For our collected hour, the $42^{nd}$ minute is slightly less active ($\sim 97.3\%$) than the average of the other $59$ minutes. Consequently, we used this correction factor to calculate the estimated number of videos per hour. The result can be seen in Fig. 1(c) and shows an interesting two-peaked temporal pattern. These estimated hours aggregate to 269.3 million video posts over a 24-hour time period.
@@
We note that the time distribution of videos per hour looks markedly different from Corso et al. [14]. We see that there are 28.5% more videos in the first minute than in the average of the other minutes, implying scheduled posts.
Of videos that are still available when we collected data, and therefore that we have metadata for, we see that there is a 14% higher posting volume at the 0th second of each minute, implying scheduled posts, or bot activity. View counts for videos posted at the zeroth second get significantly more views (p-value: 7e-5), with an average of 3,899 views at the zeroth second vs. 2,502 views for all other seconds. Videos at the 0th second more frequently use hashtags implying the desire to reach a broader audience, such as fyp, implying that these posters are using post scheduling and hashtag use to attempt to reach a broader audience.
Understanding these characteristic statistics of the platform gives very valuable priors for future work on TikTok, and we hope they serve as such for the community.
2.1.2 Engagement distributions
(a) Distribution of comment count.
(b) Distribution of view count.
(c) Distribution of like count.
(d) Distribution of share count.
(e) Distribution of video duration.
Figure 2: Histograms of various statistics in the video metadata, across the 1 hour dataset. We speculate that the peaks we see at round numbers on due to interval view count tracking optimization. However, they could also be evidence of inauthentic engagement.
@@ 2.1.3 Understanding global usage
Figure 3: Choropleth of video posting per capita for all countries, determined via the locationCreated tag available on the video post metadata. We normalize by the population of each country to show posts per capita, which specifically accounts for population rather than citizens [22]. The counts are corrected for geographical regions that we estimate that we underrepresent.
Figure 4: Regression of country population against video count in this dataset, with the 6 countries furthest from the trend line labelled, as determined by highest and lowest residuals from the linear regression. Red points correspond to the country name closest to them.
We then looked at global usage of TikTok. Research on TikTok frequently looks at country-specific slices (e.g., [23]), so understanding actual country-level usage is critical for context [20]. We infer post-country-creation from the locationCreated tag that refers to the video’s origin country  [^3] https://developers.tiktok.com/doc/research-api-specs-query-videos. While this indicator of country origin is likely confounded by VPN usage and other factors, it represents our current best method for assessing a post’s country of origin. As detailed later in Section 4.1, our method has slight variation in regional representativeness, so we adjust counts based on known statistics from our method. We explain precisely how we do this in Section 4.3.
@@
Using the locationCreated attribute, we can also infer the distribution of when TikTok posts are created in the local time, see Figure 5. In the case of countries spanning multiple timezones, we use the average hour from each timezone associated with the country from the Olson timezone database  [^5] https://github.com/stub42/pytz/blob/master/tz/zone.tab. Inaccuracies are inevitable due to lack of sub-country location information, and countries spanning multiple timezones, but the general trends should be preserved nonetheless.
@@ 2.2 Video Media
TikTok is a video-based platform, and any understanding of the nature of the platform as a whole needs to come with a look at the actual video media data of the platform.
For each post in our dataset that has associated video media data (i.e. not a story-post), we downloaded the video media data. Seeing as we collected video media content after post collection on a rolling basis, we were not able to collect media content from posts that had been deleted in the intervening time.
We wanted to assess the kind of topics discussed on the platform. We, therefore, used a combination of topic modelling and vision language models to cluster the videos and describe each cluster. We plot the posts and clusters in Figure 6.
@@
This analysis provides an exploratory glimpse at the range of topics touched on by TikTok content. We can see the broad array of these topics, with the largest being videos of a person talking directly to the camera, along with common activities in everyday life such as sports, transport, and lifestyle videos. In smaller clusters we find more uncommon topics, such as ‘Political Commentary on Police Operations in Pakistan’ (0.025% of videos), ‘Camouflaged Military Soldier Image’ (0.06%), and ‘Palestinian Flag Protest Imagery’ (0.008%).
The topic analysis completed here is not meant to be comprehensive. But by releasing the unsupervised topics discovered, we provide researchers an entrypoint to contemplating the study and engagement of particular domains within TikTok.
@@ 2.2.1 Presence of Children
The heavy use of TikTok by children and teenagers has caused concern for safety on the platform [9]. TikTok reports removing 445k accounts for child sexual exploitation and abuse  [^6] https://www.tiktok.com/transparency/en-us/child-exploitation over the first 6 months of 2024. But the lack of a representative sample of TikTok has meant so far we haven’t been able to estimate the true magnitude of child activity or the particular forms of child activity that are disproportionate on the platform. We know there is a high prevalence of children on the platform, with 50% of 5-7 year olds, and 80% of 12-15 year olds in the UK using the platform [21], but we don’t know what percentage of content on the platform features children. We aim to address that here.
We trained a classifier to detect if a video contains a child, with a mean F1 of 0.86 from 5-fold cross-validation on our train/testset (precision 0.88, recall 0.81). We provide data, model selection, and training details in Section 4.4.2. We used this classifier to find that 18.13% of the videos in our 24 hour dataset contain children (95% confidence interval: [18.06%, 18.21%] from bootstrapped confidence intervals). This is a strikingly high statistic, and underscores how deeply embedded children are in platform activity. To the best of our knowledge, no comparable large-scale studies exist for TikTok or other platforms that would allow direct benchmarking or comparison of this figure.
@@
Figure 8: Scatter plot of countries, for each country showing the share of the population under the age of 14 against the share of videos from that country containing one or more children, as determined by our video classifier. We have removed countries where there was insufficient data to make a conclusive estimate of the share of videos containing children, based on a binomial test (p=0.05). We label the top 2 and bottom 2 countries with the highest residual from the linear regression, effectively showing countries with the most overrepresented and underrepresented children in TikTok posts, respectively. Red points correspond to the country name closest to them.
We observe variability in the fraction of videos containing children by country. Figure 7 shows a choropleth of share of videos containing one or more children per country, adjusted by the share of the population under the age of 14  [^7] https://population.un.org/wpp/ (to avoid showing simply a map of countries with a youthful population). We removed countries where there were insufficient post counts to make a conclusive estimate of the share of videos containing children, based on a binomial test (p=0.05). This effectively shows countries where there is a culture of posting videos containing children. Figure 8 shows a scatter plot of countries comparing the share of the population under 14, and the share of videos containing one or more children - a linear trend is evident. We label the top 2 and bottom 2 countries with the highest residual from the linear regression, effectively showing countries with the most overrepresented and underrepresented children in TikTok posts, respectively. The variation we observe may reflect broader differences in cultural norms around public sharing of children’s images, the popularity of family-centered content, differences in platform access and usage patterns, and differences in the implementation of child protection policies across countries.
@@ 2.3 Comments
Comments are an abundant form of mass engagement data on TikTok, providing language, semantic, and temporal data. Analysis of comments provides a rich lens through which we can understand TikTok’s user base.
We collected 18.4 million comments on the videos (excluding comments on comments), associated with 13.5 million commenters. We collected these comments after the collection process was complete, in September 2024, 5 months after the day the videos were posted, using Ensemble Data  [^8] https://ensembledata.com/. The delayed comment collection process meant that we were not able to collect comments from posts that had been deleted since collection.
@@
When looking at the distribution of comment creation time depending on second of the minute and minute of the hour, we see a flat distribution (std dev. across each second of a minute: 652, std dev. across each minute of an hour: 3.7k, with a mean of 306k), indicating minimal scheduled posting of comments, and if automated comment posting is occurring, timing of posting is more sophisticated.
We also looked at average offset between video post and comment time, as shown in Figure 9. We see that for all commenters, the modal time to comment is just under 1 day, and that comments from more frequent commenters ($\geq 5$ comments in our dataset) have a modal response delay of 1 hour from post create time. We can also see that around 2/3 months after a video is posted, the commenting frequency on the video rapidly declines. This data shows us that while distribution processes on TikTok are fast, users are not engaging with the content until at least a few minutes past post time.
Looking at comment language, 47% of comments are unknown (46% of comments consist only of emojis, explaining this high percentage), 15% are in English, 9% are in Spanish, and 5% are in Arabic. The high percentage of emoji-only commenting on TikTok could imply large quantities of inauthentic activity, as emojis would be an easy way to boost a video’s engagement while requiring limited effort on behalf of the bot.
@@ 2.4 Error Analysis
A key feature of our collection method means that we receive a great number of error messages when we attempt to collect posts that either don’t exist, or are not available for other reasons. To our knowledge, this has not previously been reported on and we found it informative of internal TikTok content management practices.
We received several different types of error message for a requested video, and we report the most common error types, excluding the error message ‘item doesn’t exist’, in Table 2. We can see that an extremely large number of video posts are deleted, or privatized (status_self_see). We also gain insight into what seems to be the TikTok moderation process, with $\sim$7of videos unavailable at the time of collection due to a review stage (status_reviewing), $\sim$1of videos hidden after audit (status_audit_not_pass), and a small number potentially currently undergoing content classification for moderation (content_classification). We also see two error messages potentially related to geo-location, with cross_border_violation likely related to location-based content restrictions, for example when TikTok made new posts unavailable in Russia at the beginning of the 2022 invasion of Ukraine [27].
@@
status_deleted 3,763,773 23.21
status_self_see 3,732,155 23.01
status_reviewing 1,201,251 7.41
status_audit_not_pass 191,690 1.18
content_classification 26,654 0.16
@@
Table 2: Distribution of most common post error messages, excluding ‘item doesn’t exist’. Reported percentage is out of all videos that at some point existed in our combined datasets. So from this we can infer that around 23% of videos on TikTok are deleted. Posts can have multiple error status messages, but there is no overlap between posts with status_self_see and status_deleted, so their similar count numbers seems to be a coincidence.
When we look at deletion rates over different parts of our collection period, we see that of the posts that at some point existed (and either still exist, or taken down as evidenced by the error code), at 26 days after post time, deleted posts accounted for 12% of posts, and at 215 days after post time, they accounted for 20% of posts. TikTok reports only removing 1% of posts  [^9] https://www.tiktok.com/transparency/en-us/community-guidelines-enforcement-2024-4, indicating that these are largely user removals.
This data gives us a first insight into inaccessible content on TikTok, and future work can leverage these statistics to bring to light the moderation processes used by the platform, the rate of deletion, and the nature of cross-border content restrictions.